The second-order rate constant for the following gas-phase reaction is 0.043 1/MLaTeX: \cdot?s. We start with 0.497 mol C2F4 in a 2.23 liter container, with no C4F8 initially present.   C2F4  LaTeX: \longrightarrow? 1/2 C4F8

What will be the concentration of C4F8 after 2.17 hours? Enter to 4 decimal places. HINT: You will need to figure out how much of C2F4 was used (that is the only material that can be converted to the product).
Answer #1

For second order reaction

= (1/A) = 1/(A°) + k*T

= Time = 2.17*60*60 = 7812 sec.

Now molarity of C2F4 = moles / volume

= 0.497/2.23 = 0.223 M

= 1/A = (1/0.223) + 0.043*7812

= A = 2.94 *10^-3 M

= This amount is finally remaining

= Change in amount of C2F4 = (0.497)-(2.94*10^-3)

= 0.49406

= Balance reaction says that 1 mole convert into 0.5 mole of product

= Moles of C4F8 = (0.49406) / 2 = 0.24703 M

If you satisfied with the solution please like it thankxx